# DualNILM: Energy Injection Identification Enabled Disaggregation with Deep Multi-Task Learning

### v1

@@ VII Conclusion
We presented DualNILM, a deep multi-task learning framework tailored to address the challenges of NILM in the presence of behind-the-meter energy injection, such as micro-inverters. By jointly performing appliance state recognition and energy injection disaggregation within a Transformer-based architecture, DualNILM effectively captures multi-scale temporal dependencies and leverages the interrelated nature of these tasks. Extensive experiments on real-world laboratory datasets with micro-inverter injection, as well as synthesized PV injection datasets derived from well-known NILM benchmarks, demonstrate that DualNILM significantly outperforms both conventional and deep learning-based methods, maintaining robust performance even under challenging conditions.
Despite these promising results, applying DualNILM to more complex household scenarios, such as those involving intricate appliance interactions (e.g., Vehicle-to-Grid systems [49] or home battery management systems [50, 51]), remains a challenge. However, the flexibility of the DualNILM framework provides a strong foundation for future improvements. Incorporating additional features, such as environmental data, or integrating contextual insights from large language models (LLMs) [52, 53], presents opportunities to enhance its performance and extend its applicability. The inclusion of synthesized PV injection datasets, which approximate real-world conditions, also serves as a valuable resource for advancing NILM research. These datasets enable researchers to investigate NILM in the context of renewable energy integration and uncover deeper challenges in this domain.
In conclusion, our work demonstrates strong potential for advancing NILM solutions in modern energy systems with Behind-the-Meter energy injection. Addressing the aforementioned challenges is crucial for developing smarter, more sustainable energy management systems in the era of increasing renewable energy adoption.
Appendix A Theoritical Analysis of the Challenge: When Negative Flows Disrupt Classical NILM
In the classic NILM framework, the observed signal (where is the number of time samples) is seen as a single-channel mixture of multiple component signals (appliances), each assumed to be nonnegative over time. When behind-the-meter injection is introduced, at least one component can take negative values, increasing the dimensionality of the solution space and weakening certain identifiability properties that hold under nonnegativity constraints. We make this more precise below.
@@ A-A Matrix Formulation of NILM
Consider appliances in a building (including the “injection appliance”). For each appliance , define
where is the instantaneous power consumption (or production, if negative). A binary activation sequence
indicates whether appliance is operating at each time . Thus, the aggregate measured load satisfies
@@
A-B Relaxation to Negative Flows and Identifiability
When behind-the-meter solar (or other storage/generation) is present, the corresponding for that source may take negative values at certain time points—i.e.,
Hence, at least one column in the set is no longer restricted to but spans , effectively dilating the feasible solution set:
In words, once any is free to be negative, the dimension of this feasible set grows, potentially yielding multiple plausible explanations of the same aggregate signal .
Loss of Uniqueness and Identifiability.
Under classical NILM assumptions (fully nonnegative), unique or near-unique decompositions can sometimes be proven when the underlying sources have distinct support or exhibit certain “sparse plus low-rank” structures. These uniqueness arguments are often reminiscent of theorems in blind source separation (BSS) and NMF, where the nonnegative cone enforces a geometric structure that “pins down” solutions. Once negative values are allowed (even for a single source), the problem is closer to an unconstrained BSS, for which uniqueness is well known to degrade. In more formal terms, NMF-based identifiability relies on the fact that data points lie in a convex cone generated by the basis columns [29]; introducing negative components breaks those convex cone arguments.
A-C Signal Space Interpretation and Rank Constraints
@@
where represents a small dictionary of “candidate appliance patterns” for device , and are nonnegative coefficients that specify which pattern from is used at each time. In a purely consumptive (classical) case, imposes a conic constraint. For behind-the-meter injection, the dictionary for the injecting device must allow negative entries. Thus, while the classical (nonnegative) scenario yields
the new scenario broadens the feasible set to
@@
In short, allowing negative flows expands the solution space beyond the convex cone imposed by nonnegative appliances. This expansion degrades the ability to isolate unique solutions and increases the difficulty of load disaggregation.
@@ B-A Model Hyperparameters
The overall architecture of DualNILM is illustrated in Fig. 2. Our model comprises four main components: 1) CNN Encoders, 2) Transformer Encoders, 3) a Transformer Decoder, and 4) Task-specific projection layers. This design allows for the effective capture of both local and global temporal dependencies in the power consumption patterns while enabling multi-task learning.
@@
Appendix C Datasets and Solar Simulation Details
@@ C-A1 Data Alignment and Resampling
C-A2 Appliance State Labeling Thresholds
@@
C-A3 Data Splitting
The datasets are split into training and testing sets based on the dates specified in Section V-A2.
@@ C-A3 Data Splitting
For cross-validation purposes, we implement three distinct evaluation modes: (1) standard chronological split as specified above, (2) testing on the initial period while training on the remainder, and (3) testing on a mid-period segment while training on the surrounding data. All test periods maintain consistent durations across different splits for fair comparison.
C-A4 Simulation Details
Our data processing pipeline is implemented using Python and leverages libraries such as pandas for data manipulation and NumPy for numerical computations. The code for the simulation and data processing will be available after this paper is accepted.
@@ Appendix D Details of Experiment
The dataset and code used in this study will be made publicly available after acceptance, providing a benchmark for future research in NILM with energy injection.
@@ D-A Details of Benchmark Methods
Appendix E More Details of Experiment Result
@@
E-A Overview of all Results Cross all Dataset
TableVII and VIII show the overall average performance metrics on appliance state recognization and energy injection disaggregation tasks.
@@ E-A Overview of all Results Cross all Dataset
| Laboratory Dataset (Real Behind-the-Meter Injection) | FHMM | 60.95 | 28.32 | 16.84 | 17.25 | |||
@@
| Laboratory Dataset (Real Behind-the-Meter Injection) | FHMM | 0.5015 | 0.4492 | |||||
@@ E-B Laboratory Dataset Results
Table. IX shows the details of the laboratory dataset performance metrics for DualNILM and benchmarks.
Fig. 5 shows all seven different appliances (including micro-inverter)’s ground truth state, DualNILM, and benchmarks prediction on Dec.8, 2023 of our dataset.

### v2

@@ VII Conclusion
We presented DualNILM, a deep multi-task learning framework tailored to address the challenges of NILM in the presence of BTM energy injection, such as micro-inverters and photovoltaic system outputs. Extensive experiments on real-world laboratory datasets with micro-inverter injection, as well as synthesized PV injection datasets derived from well-known NILM benchmarks, demonstrate DualNILM’s superior performance and the strong potential for advancing NILM solutions in modern energy systems with BTM energy injection.
Our findings underscore DualNILM’s potential for robust NILM in modern energy systems that integrate renewable energy sources, and our provided open research infrastructure will help the NILM community continue to advance research in this critical direction.
Appendix A Theoretical Analysis of the Challenge: When Negative Flows Disrupt Classical NILM
In the classic NILM framework, the observed signal (where is the number of time samples) is seen as a single-channel mixture of multiple component signals (appliances), each assumed to be nonnegative over time. When BTM injection is introduced, at least one component can take negative values, increasing the dimensionality of the solution space and weakening certain identifiability properties that hold under nonnegativity constraints.
@@ A-A Matrix Formulation of NILM
Consider appliances in a building (including the ”injection appliance”). For each appliance , define:
where is the instantaneous power consumption (or production, if negative). A binary activation sequence indicates whether appliance is operating at each time . Thus, the aggregate measured load satisfies:
@@
A-B Why BTM Injection Complicates NILM
Weakened Identifiability: Nonnegative factorizations often depend on convexity-based geometric constraints; negative entries undermine these constraints and can make standard disaggregation fall back on naive assumptions.
A-C Relaxation to Negative Flows and Identifiability
When behind-the-meter solar (or other storage/generation) is present, the corresponding for that source may take negative values at certain time points:
Hence, at least one column in the set is no longer restricted to but spans , effectively dilating the feasible solution set as shown in Equation 21.
Degraded Identifiability and Solution Stability. Under classical NILM assumptions (fully nonnegative), more stable decompositions can sometimes be achieved when the underlying sources have distinct support or exhibit certain ”sparse plus low-rank” structures. These identifiability arguments are often reminiscent of theorems in blind source separation (BSS) and NMF, where the nonnegative cone enforces a geometric structure that ”pins down” solutions. Once negative values are allowed (even for a single source), the problem becomes closer to an unconstrained BSS, for which solution identifiability is well known to deteriorate significantly. In more formal terms, NMF-based identifiability relies on the fact that data points lie in a convex cone generated by the basis columns [45]; introducing negative components breaks those convex cone arguments.
A-D Signal Space Interpretation and Rank Constraints
@@
where represents a small dictionary of ”candidate appliance patterns” for device , and are nonnegative coefficients that specify which pattern from is used at each time.
In a purely consumptive (classical) case, imposes a conic constraint. For BTM injection, the dictionary for the injecting device must allow negative entries. Thus, while the classical (nonnegative) scenario yields:
the new scenario broadens the feasible set to:
@@
In summary, allowing negative flows expands the solution space beyond the convex cone imposed by nonnegative appliances. This expansion degrades the ability to achieve stable solution identification and significantly increases the computational complexity of load disaggregation. Traditional NILM methods that treat all loads as nonnegative typically fail to disentangle consumption from behind-the-meter generation when measurements dip below zero or remain subdued for extended intervals, as demonstrated in our experimental results (Figure 1 and Table III).
@@ B-A Model Hyperparameters
The overall architecture of DualNILM is illustrated in Figure 2. Our model comprises four main components: 1) CNN Encoders, 2) Transformer Encoders, 3) a Transformer Decoder, and 4) Task-specific projection layers. This design allows for the effective capture of both local and global temporal dependencies in the power consumption patterns while enabling multi-task learning.
@@
Appendix C Datasets and PV Simulation Details
@@ C-A1 Data Alignment and Resampling
C-A2 The Choice of Simulated 2kW PV Scale
For the context of REDD and UKDALE, in the early-2010s PV market, residential new-construction PV systems most commonly fell in the 2-4 kW range [46], which puts our choice of 2 kW injection at the small-system end but still representative for that period.
From a siting perspective, NREL’s 2010 residential benchmark assumes 35 m2 of array area using 27 modules (14.5% efficiency) for a 5 kWdc system; by simple proportionality, a 2 kW array needs 14 m2, well within the envelope of typical roof planes [47].
Consistently, NREL’s rooftop technical-potential study requires just one contiguous roof plane of 10 m2, noting this supports 1.5 kW at 15% efficiency, again indicating that 2 kW is feasible on modest roof areas [48].
Policy design at the time also discouraged oversizing: California’s NSHP required systems to be “sized so that the electricity produced offsets part or all of on-site needs,” with a minimum of 1 kW AC and systems 7.5 kW AC deemed “sized to serve on-site load.” This aligns with our decision not to exceed historical maximum household loads in these datasets [49].
C-A3 Appliance State Labeling Thresholds
@@
C-A4 Data Splitting
The datasets are split into training and testing sets based on the dates specified in Section V-A.
@@ C-A4 Data Splitting
For cross-validation purposes, we implement three distinct evaluation modes: (1) standard chronological split as specified above, (2) testing on the initial period while training on the remainder, and (3) testing on a mid-period segment while training on the surrounding data. For fair comparison, all test periods maintain consistent durations across different splits.
C-A5 Simulation Details
Our data processing pipeline is implemented using Python and leverages libraries such as pandas for data manipulation and NumPy for numerical computations.
@@ D-A Details of Benchmark Methods
Appendix E Full Experiment Results
@@
E-A Overview of All Results Cross All Datasets
Table VII and VIII show the overall average performance metrics on appliance state recognition and energy injection disaggregation tasks.
@@ E-A Overview of All Results Cross All Datasets
| Laboratory Dataset (Real BTM Injection) | FHMM | 60.95 | 28.32 | 16.84 | 17.25 | |||
@@
| Laboratory Dataset (Real BTM Injection) | FHMM | 0.5015 | 0.4492 | |||||
@@ E-B Laboratory Dataset Results
Table IX shows the details of the laboratory dataset performance metrics for DualNILM and benchmarks.
Figure 5 shows all seven different appliances (including micro-inverter)’s ground truth state, DualNILM, and benchmarks prediction on Dec.8, 2023 of our dataset.
@@ E-D3 Overall on UKDALE
Appendix F Discussion of Limitation & Future Works
Despite these promising results, applying DualNILM to more complex household scenarios, such as those involving intricate appliance interactions (e.g., Vehicle-to-Grid systems [50] or home battery management systems [51, 52]), remains a challenge. However, the flexibility of the DualNILM framework provides a strong foundation for future improvements. Incorporating additional features, such as environmental data, or integrating contextual insights from large language models (LLMs) [53, 54], presents opportunities to enhance its performance and extend its applicability.
What’s more, the provided synthesized PV injection datasets, which approximate real-world conditions, also serve as a valuable resource for advancing NILM research. These datasets enable researchers to investigate NILM in the context of renewable energy integration and uncover deeper challenges in this domain.